Three times the mystery number was increased by 100, and the result was 12 less than the number. Find the mystery number.
Answer by checkley77(12844)

3x+100=x-12
3x-x=-12-100
2x=-112
x=-112/2
x=-56 answer.
proof
3*-56+100=-56-12
-168+100=-68
-68=-68